# Uniform Hyperbolicity and Symbolic Dynamics: Markov Partitions, Shadowing, and the Coding of Axiom A Systems

### 19 Apr 2026 v1

@@ Abstract.
This Part establishes the geometric theory of uniformly hyperbolic sets with explicit quantitative bounds throughout, and contains five main theorems. The Stable Manifold Theorem is proved via the backward graph transform, with a complete fiber-contraction argument yielding $C^{r}$ regularity and Hölder dependence of the local stable and unstable manifolds on the base point, with explicit manifold-size estimate in terms of the contraction rate $\lambda$ and the second-derivative bound of the diffeomorphism. The Spectral Decomposition Theorem gives the unique decomposition of the nonwandering set into basic sets, with explicit mixing rates for the topologically mixing factors. The Shadowing Lemma provides explicit error bounds controlling how far a pseudo-orbit deviates from a tracking true orbit. The existence of Markov partitions of arbitrarily small diameter is established constructively, with explicit diameter bounds expressed in terms of the shadowing constants. Finally, the coding map from the subshift of finite type to the hyperbolic set is constructed and shown to be Hölder continuous with quantitative control on the exceptional set where it fails to be injective. Along the way we establish canonical coordinates through the bracket map with quantitative bounds. All constants are expressed in terms of the contraction rate, the Hölder exponent of the derivative, the manifold dimension, and the injectivity radius, providing the quantitative infrastructure required to transfer the symbolic spectral theory of Part I and the variational theory of Part II to the smooth setting. This Part constitutes Part III of a six-part series on the thermodynamic formalism for hyperbolic dynamical systems.
@@
Fields Medalist, Professor at the Collège de France,
who introduced the author to hyperbolic dynamics.
@@ 1. Introduction
The geometric theory of hyperbolic dynamical systems originates in the work of Hadamard (1898) on geodesics on negatively curved surfaces, Morse (1921) on symbolic coding of geodesic flows, and the foundational contributions of Anosov (1967) and Anosov and Sinai (1967) on globally hyperbolic diffeomorphisms. This Part establishes the geometric theory of Axiom A diffeomorphisms with quantitative estimates throughout, providing the bridge that connects the symbolic thermodynamic formalism of Parts I–II to smooth dynamics. The main results are: a complete proof of the Stable Manifold Theorem via the graph transform with Hölder exponent $\alpha=\log\mu/\log\lambda$ for the dependence on the base point (this exponent is known from Hasselblatt (1994); we compute the Hölder constant explicitly); canonical coordinates with quantitative bracket map estimates; the Spectral Decomposition Theorem with explicit mixing rates; a shadowing lemma with explicit constants; Markov partitions with diameter bounds in terms of hyperbolicity data; and the coding map $\pi:\Sigma_{A}\to\Lambda$ with quantitative control on the exceptional set.
@@
The coding map $\pi:\Sigma_{A}\to\Lambda$ constructed in Section 9 satisfies three properties required by Parts IV–VI: it intertwines the shift $\sigma$ and the diffeomorphism $f$, it is Hölder continuous with explicit exponent, and it is injective off a set of measure zero for any Gibbs measure. Given a Hölder potential $\phi:\Lambda\to\mathbb{R}$, the composition $\phi\circ\pi$ is Hölder on $\Sigma_{A}$, and the equilibrium state for $\phi$ on $\Lambda$ is the pushforward under $\pi$ of the equilibrium state for $\phi\circ\pi$ on $\Sigma_{A}$, which exists and is unique by the theory developed in Part I Thiam (2025a).
@@
This Part is the third of a six-part series on the thermodynamic formalism for hyperbolic dynamical systems. Part I Thiam (2025a) establishes the spectral theory of Gibbs measures on symbolic spaces, building on the transfer operator methods of Baladi (2000) and Parry and Pollicott (1990), the $g$-measure theory of Keane (1972), the conformal measures of Denker and Urbański (1991), and the perturbation theory of Kato (1980). Part II Thiam (2025b) develops the convex-analytic structure of the pressure functional. The present Part provides the geometric bridge. Parts IV–VI develop the transfer operator on manifolds, statistical limit theorems, and structural consequences. The extension to partial hyperbolicity was initiated by Brin and Pesin (1974) and developed by Pugh and Shub (1997, 2000); see Bonatti et al. (2005) and Crovisier and Potrie (2015) for modern treatments. Mañéś ergodic closing lemma Mañé (1982) provides complementary orbit approximation tools. The thermodynamic formalism connecting hyperbolic dynamics to statistical mechanics was developed by Sinai (1972), Ruelle (1968, 1973, 1976, 1978), and Bowen (1970a, 1974); Bowen and Ruelle (1975) extended the theory to Axiom A flows. Beyond uniform hyperbolicity, Young (1998) developed statistical properties via tower constructions (see Young (2002) for a survey of SRB measures), Climenhaga and Thompson (2012) proved intrinsic ergodicity for $\beta$-shifts and their factors, established uniqueness of equilibrium states via obstruction entropies Climenhaga and Thompson (2014), and extended these results to flows and homeomorphisms with non-uniform hyperbolic structure Climenhaga and Thompson (2016). de la Llave (1997) developed smooth conjugacy theory, and Berghout et al. (2019) clarified the relationship between Gibbs and $g$-measures.
@@
The non-uniformly expanding extension of the Markov partition and coding program of our Part III was developed by Luzzatto and Tucker (1999), who constructed induced Markov structures for maps with criticalities and singularities and established decay of return times in Publications Mathématiques de l’IHÉS. The comprehensive survey of Luzzatto (2006) in the Handbook of Dynamical Systems collects the methods and results of this line of work. Alves et al. (2005) constructed the Markov tower in the non-uniformly expanding setting and established polynomial and exponential rates of decay of correlations depending on the decay rate of the return times. Alves et al. (2000) proved existence of SRB measures for partially hyperbolic systems whose central direction is mostly expanding, a setting closely adjacent to ours. Our Part III is the uniformly hyperbolic anchor: we construct Markov partitions with explicit diameter bounds, controlled by the contraction rate and the injectivity radius, and prove the coding map is Hölder continuous with quantitative bounds on the exceptional set. The works cited above do not provide such explicit constants.
@@
This Part is organized as follows. Section 2 fixes notation and collects the prerequisites from Riemannian geometry, function spaces, spectral theory, measure theory, and symbolic dynamics that are used in the sequel. Section 3 develops the quantitative theory of hyperbolic sets: the existence of adapted metrics with explicit contraction rate $\lambda$, the Hölder continuity of the stable-unstable splitting with exponent computed from the hyperbolicity data, and the basic properties of Axiom A diffeomorphisms and their periodic orbits. Section 4 proves Main Theorem 4.2 (the Stable Manifold Theorem) via the backward graph transform, establishing the existence, $C^{r}$ regularity, exponential contraction, and Hölder dependence on base points of the local stable and unstable manifolds, with explicit size estimate $\varepsilon_{0}=(1-\lambda)^{2}/(4C_{0})$. Section 5 establishes canonical coordinates through the bracket map $[\cdot,\cdot]$, proves the local product structure of hyperbolic sets, and quantifies expansiveness and the fundamental neighborhood; these are the geometric tools used in the Markov partition construction. Section 6 proves Main Theorem 6.1 (the Spectral Decomposition), which partitions the nonwandering set $\Omega(f)$ into finitely many closed $f$-invariant basic sets on which $f$ is topologically transitive, and establishes the cyclic refinement into mixing components along with quantitative mixing rates. Section 7 proves Main Theorem 7.3 (the Shadowing Lemma) with explicit constants, derives the Anosov Closing Lemma and the specification property as consequences, and develops quantitative periodic orbit approximation estimates needed in Parts IV–VI. Section 8 proves Main Theorem 8.5 (Existence of Markov Partitions), constructing Markov partitions of arbitrarily small diameter on each basic set via the shadowing lemma, with explicit diameter bounds in terms of the hyperbolicity data. Section 9 proves Main Theorem 9.8 (Symbolic Coding), establishing the Hölder continuous coding map $\pi:\Sigma_{A}\to\Omega$ with quantitative control on the exceptional set where $\pi$ fails to be injective, and identifies the topological entropy of $f|_{\Omega}$ with $\log\rho(A)$. The dimension theory of hyperbolic sets, developed by Pesin (1977, 1997), Young (1982), and Barreira et al. (1999), relies on the coding map and Markov partition constructed here. Section 10 provides a complete numerical illustration for the Smale horseshoe, computing explicit values of the adapted metric constant, stable manifold size, shadowing constant, Markov partition diameter, and grid resolution needed for rigorous symbolic coding at prescribed accuracy. Section 11 concludes the Part with a summary of the five Main Theorems and open problems. The appendix collects technical proofs and supplementary material: prerequisite results on Hölder spaces and quasi-compact operators, and standard measure-theoretic lemmas used throughout.
@@ A $C^{r}$ diffeomorphism $f:M\to M$ satisfies Axiom A if:
For an Anosov diffeomorphism, $\Omega(f)=M$ (every point is nonwandering since the stable and unstable manifolds of any point are dense). The density of periodic points follows from the Anosov Closing Lemma (Theorem 7.5), proved in Section 7. See Corollary 7.6. ∎
@@ 4.5. Quantitative Estimates
Step 3: Sharpening to rate $\lambda$. The bound in Step 2 has rate $\lambda^{\prime}=(5\lambda+3)/8$, not $\lambda$. To obtain rate $\lambda$ as stated, we use that $\lambda$ is the hyperbolicity exponent of the adapted metric, chosen with a safety factor in Section 4.2. Concretely, the adapted metric construction (Theorem 3.5) gives a metric $\|\cdot\|_{*}$ in which $\|Df|_{E^{s}}\|_{*}\leq\lambda$. Choosing the adapted metric with rate $\lambda_{0}<\lambda$ (any rate strictly less than $\lambda$ is allowed as long as $\lambda_{0}$ exceeds the true contraction rate of $Df|_{E^{s}}$), the same calculation as above yields rate $\lambda_{0}+C_{0}\varepsilon(1+K)$, which for $\varepsilon$ sufficiently small is at most $\lambda$. Rescaling constants accordingly proves the claim with rate exactly $\lambda$. ∎

### 23 Apr 2026 v2

# Uniform Hyperbolicity and Symbolic Dynamics: Markov Partitions, Shadowing, and the Coding of Axiom A Diffeomorphisms

@@ Abstract.
This Part establishes the geometric theory of uniformly hyperbolic sets with explicit quantitative bounds throughout, and contains five main theorems. The Stable Manifold Theorem is proved via the backward graph transform, with a complete fiber-contraction argument yielding $C^{r}$ regularity and Hölder dependence of the local stable and unstable manifolds on the base point, with explicit manifold-size estimate in terms of the contraction rate $\lambda$ and the second-derivative bound of the diffeomorphism. The Spectral Decomposition Theorem gives the unique decomposition of the nonwandering set into basic sets, with explicit mixing rates for the topologically mixing factors. The Shadowing Lemma provides explicit error bounds controlling how far a pseudo-orbit deviates from a tracking true orbit. The existence of Markov partitions of arbitrarily small diameter is established constructively, with explicit diameter bounds expressed in terms of the shadowing constants. Finally, the coding map from the subshift of finite type to the hyperbolic set is constructed and shown to be Hölder continuous with quantitative control on the exceptional set where it fails to be injective. Along the way we establish canonical coordinates through the bracket map with quantitative bounds. All constants are expressed in terms of the contraction rate, the Hölder exponent of the derivative, the manifold dimension, and the injectivity radius, providing the quantitative infrastructure required to transfer the symbolic spectral theory of Part I Thiam (2026a) and the variational theory of Part II Thiam (2026b) to the smooth setting. This Part constitutes Part III of a six-part series on the thermodynamic formalism for hyperbolic dynamical systems.
@@
Fields Medalist and Professor at the Collège de France, with whom the author
had the privilege of working, and who introduced him to hyperbolic dynamics.
@@ 1. Introduction
The geometric theory of hyperbolic dynamical systems originates in the work of Hadamard (1898) on geodesics on negatively curved surfaces, Morse (1921) on symbolic coding of geodesic flows, and the foundational contributions of Anosov (1967) and Anosov and Sinai (1967) on globally hyperbolic diffeomorphisms. This Part establishes the geometric theory of Axiom A diffeomorphisms with quantitative estimates throughout, providing the bridge that connects the symbolic thermodynamic formalism of Parts I–II Thiam (2026a, b) to smooth dynamics. The main results are: a complete proof of the Stable Manifold Theorem via the graph transform with Hölder exponent $\alpha=\log\mu/\log\lambda$ for the dependence on the base point (this exponent is known from Hasselblatt (1994); we compute the Hölder constant explicitly); canonical coordinates with quantitative bracket map estimates; the Spectral Decomposition Theorem with explicit mixing rates; a shadowing lemma with explicit constants; Markov partitions with diameter bounds in terms of hyperbolicity data; and the coding map $\pi:\Sigma_{A}\to\Lambda$ with quantitative control on the exceptional set.
@@
The coding map $\pi:\Sigma_{A}\to\Lambda$ constructed in Section 9 satisfies three properties required by Parts IV–VI Thiam (2026d, e, f): it intertwines the shift $\sigma$ and the diffeomorphism $f$, it is Hölder continuous with explicit exponent, and it is injective off a set of measure zero for any Gibbs measure. Given a Hölder potential $\phi:\Lambda\to\mathbb{R}$, the composition $\phi\circ\pi$ is Hölder on $\Sigma_{A}$, and the equilibrium state for $\phi$ on $\Lambda$ is the pushforward under $\pi$ of the equilibrium state for $\phi\circ\pi$ on $\Sigma_{A}$, which exists and is unique by the theory developed in Part I Thiam (2026a).
@@
The contributions of this Part are fivefold, corresponding to its five Main Theorems. First, we prove the Stable Manifold Theorem (Main Theorem 4.2) with explicit manifold-size estimate $\varepsilon_{0}=(1-\lambda)^{2}/(4C_{0})$ where $C_{0}=\max(\|D^{2}f\|_{\infty},\|D^{2}f^{-1}\|_{\infty})$, $C^{r}$ regularity obtained through the fiber-contraction principle applied to successive derivatives of the graph map, exponential contraction $d(f^{n}(x),f^{n}(y))\leq\lambda^{n}d(x,y)$ along the manifolds, and Hölder dependence of the local stable manifold on the base point with explicit exponent $\alpha_{s}=\beta\log\lambda^{-1}/(\log\|Df\|_{\infty}+\beta\log\lambda^{-1})$ for $f\in C^{1+\beta}$. Second, we prove the Spectral Decomposition Theorem (Main Theorem 6.1) giving the unique decomposition $\Omega(f)=\Omega_{1}\cup\cdots\cup\Omega_{s}$ of the nonwandering set into finitely many closed $f$-invariant basic sets on which $f$ is topologically transitive, together with the cyclic refinement $\Omega_{i}=X_{1,i}\cup\cdots\cup X_{n_{i},i}$ into topologically mixing components. Third, we prove the Shadowing Lemma (Main Theorem 7.3) with explicit linear dependence $\alpha=C^{-1}(1-\lambda)\beta$ between the pseudo-orbit tolerance $\alpha$ and the shadowing accuracy $\beta$, where the constant $C$ depends on the local product structure scale and the adapted metric distortion $K=c(1-\lambda/\mu)^{-1}$; the Anosov Closing Lemma and the specification property are derived as consequences with explicit constants. Fourth, we prove the Existence of Markov Partitions (Main Theorem 8.5) constructively, obtaining partitions of arbitrarily small diameter on each basic set through a refinement argument on an initial cover of shadowing points, with the diameter bound $\max_{i}\mathrm{diam}(R_{i})\leq C\cdot\beta$ expressed directly in terms of the shadowing accuracy. Fifth, we prove the Symbolic Coding Theorem (Main Theorem 9.8) establishing the coding map $\pi:\Sigma_{A}\to\Omega$ as a continuous surjection intertwining the shift with the diffeomorphism, Hölder continuous with accuracy bound $\mathrm{diam}(K_{N}(a))\leq C\lambda^{N}\mathrm{diam}(\mathcal{R})$, injective off the countable boundary set $\bigcup_{n}f^{n}(\partial\mathcal{R})$ of zero measure for every ergodic invariant measure; the topological entropy of $f|_{\Omega}$ is identified with $\log\rho(A)$. All constants are tracked explicitly as functions of the contraction rate $\lambda$, the Hölder exponent of $Df$, the manifold dimension $d$, and the injectivity radius of $M$, providing the quantitative infrastructure required by Parts IV–VI Thiam (2026d, e, f).
Our technical approach combines the graph-transform construction of invariant manifolds with the bracket-based construction of the Markov partition. The Stable Manifold Theorem is proved through the backward graph transform $\Gamma^{-1}$ on the space of Lipschitz graphs $\phi:B^{s}_{\delta}\to B^{u}_{\delta}$: since the linearization $B$ in the unstable direction satisfies $\|B^{-1}\|\leq\lambda<1$, the forward graph transform $\Gamma$ is expansive, while $\Gamma^{-1}$ is a strict contraction in the $C^{0}$ norm with explicit rate $\theta=\lambda/(1-\lambda C_{1}\delta(K+1))$. The unique fixed point of $\Gamma^{-1}$ is the graph of the local stable manifold. $C^{r}$ regularity is obtained inductively by applying the fiber-contraction principle at each order $j\leq r$: formal differentiation of the fixed-point equation yields a linear operator $\Psi_{j}$ on the space of $j$-multilinear-map-valued functions, with contraction rate bounded by $\lambda^{j+1}(1+\mathcal{O}(\delta))$, and the unique fixed point is identified with $D^{j}\phi_{\infty}$. Hölder dependence of the manifold on the base point is obtained through a truncation-and-balance argument: the forward orbit of the base-point comparison is tracked for $N\sim\log(1/d(x,y))/\log\|Df\|_{\infty}$ iterates and the telescoping error is balanced against the exponential decay of the comparison. For the geometric theory, canonical coordinates through the bracket map $[x,y]=W^{s}_{\varepsilon}(x)\cap W^{u}_{\varepsilon}(y)$ provide a local product structure on the hyperbolic set, with Lipschitz constant $C_{1}=(1-K)^{-1}$ where $K$ is the Lipschitz constant of the manifold graphs. Expansiveness follows from the local product structure: any pair of orbits staying within $\varepsilon<\delta_{0}$ of each other must coincide by $W^{s}_{\varepsilon}(x)\cap W^{u}_{\varepsilon}(x)=\{x\}$. For the Shadowing Lemma, we iterate the bracket map along the pseudo-orbit: given $\alpha$-pseudo-orbit $\{x_{i}\}$, we construct a shadowing point by successive bracketing of $f^{M}(x^{\prime}_{(k-1)M})$ with $x_{kM}$, and the telescoping estimate yields the linear bound $\alpha=C^{-1}(1-\lambda)\beta$. The Markov partition construction begins with a $\gamma$-dense set $P\subset\Omega$ of reference points, coded into a subshift $\Sigma(P)$ of finite type via the pseudo-orbit condition, and the shadowing map $\theta:\Sigma(P)\to\Omega$ defines initial rectangles $T_{s}=\theta(\{q:q_{0}=p_{s}\})$. The final Markov partition is obtained by refining these rectangles according to all four possible intersection patterns with other rectangles in the cover, eliminating boundary overlaps to produce the open rectangles $R(x)$ with the required Markov property. The coding map is then $\pi(a)=\bigcap_{n}f^{-n}(R_{a_{n}})$: successive forward preimages narrow the unstable direction by factor $\lambda^{n}$ and backward iterates narrow the stable direction by factor $\lambda^{n}$, so the intersection collapses to a single point with explicit accuracy $\mathrm{diam}(K_{N}(a))\leq C\lambda^{N}\mathrm{diam}(\mathcal{R})$. These methods replace the compactness arguments used in Bowen (1975) and the semi-conjugacy techniques of Sinai (1968a, b) by constructive, quantitative procedures.
This Part is organized as follows. Section 2 fixes notation and collects prerequisites from Riemannian geometry, function spaces, spectral theory, measure theory, and symbolic dynamics. Section 3 develops the quantitative theory of hyperbolic sets, including the adapted metric construction and the Hölder continuity of the stable-unstable splitting. Section 4 proves Main Theorem 4.2. Section 5 establishes canonical coordinates through the bracket map and proves the local product structure. Section 6 proves Main Theorem 6.1. Section 7 proves Main Theorem 7.3 and derives the Anosov Closing Lemma and the specification property as consequences. Section 8 proves Main Theorem 8.5. Section 9 proves Main Theorem 9.8 and identifies the topological entropy of $f|_{\Omega}$ with $\log\rho(A)$. The dimension theory of hyperbolic sets, developed by Pesin (1977, 1997), Young (1982), and Barreira et al. (1999), relies on the coding map and Markov partition constructed here. Section 10 provides a numerical illustration for the Smale horseshoe, computing explicit values of the adapted metric constant, stable manifold size, shadowing constant, Markov partition diameter, and grid resolution for rigorous symbolic coding at prescribed accuracy. Section 11 concludes with a summary of the five Main Theorems and open problems. The appendix collects technical proofs and supplementary material on Hölder spaces, quasi-compact operators, and measure-theoretic lemmas.
@@
The non-uniformly expanding extension of the Markov partition and coding program developed in the present Part was extended by Luzzatto and Tucker (1999), who constructed induced Markov structures for maps with criticalities and singularities and established decay of return times in Publications Mathématiques de l’IHÉS. The comprehensive survey of Luzzatto (2006) in the Handbook of Dynamical Systems collects the methods and results of this line of work. Alves et al. (2005) constructed the Markov tower in the non-uniformly expanding setting and established polynomial and exponential rates of decay of correlations depending on the decay rate of the return times. Alves et al. (2000) proved existence of SRB measures for partially hyperbolic systems whose central direction is mostly expanding, a setting closely adjacent to ours. The present Part is the uniformly hyperbolic anchor: we construct Markov partitions with explicit diameter bounds, controlled by the contraction rate and the injectivity radius, and prove the coding map is Hölder continuous with quantitative bounds on the exceptional set. The works cited above do not provide such explicit constants.
@@
This Part is organized as follows. Section 2 fixes notation and collects the prerequisites from Riemannian geometry, function spaces, spectral theory, measure theory, and symbolic dynamics that are used in the sequel. Section 3 develops the quantitative theory of hyperbolic sets: the existence of adapted metrics with explicit contraction rate $\lambda$, the Hölder continuity of the stable-unstable splitting with exponent computed from the hyperbolicity data, and the basic properties of Axiom A diffeomorphisms and their periodic orbits. Section 4 proves Main Theorem 4.2 (the Stable Manifold Theorem) via the backward graph transform, establishing the existence, $C^{r}$ regularity, exponential contraction, and Hölder dependence on base points of the local stable and unstable manifolds, with explicit size estimate $\varepsilon_{0}=(1-\lambda)^{2}/(4C_{0})$. Section 5 establishes canonical coordinates through the bracket map $[\cdot,\cdot]$, proves the local product structure of hyperbolic sets, and quantifies expansiveness and the fundamental neighborhood; these are the geometric tools used in the Markov partition construction. Section 6 proves Main Theorem 6.1 (the Spectral Decomposition), which partitions the nonwandering set $\Omega(f)$ into finitely many closed $f$-invariant basic sets on which $f$ is topologically transitive, and establishes the cyclic refinement into mixing components along with quantitative mixing rates. Section 7 proves Main Theorem 7.3 (the Shadowing Lemma) with explicit constants, derives the Anosov Closing Lemma and the specification property as consequences, and develops quantitative periodic orbit approximation estimates needed in Parts IV–VI Thiam (2026d, e, f). Section 8 proves Main Theorem 8.5 (Existence of Markov Partitions), constructing Markov partitions of arbitrarily small diameter on each basic set via the shadowing lemma, with explicit diameter bounds in terms of the hyperbolicity data. Section 9 proves Main Theorem 9.8 (Symbolic Coding), establishing the Hölder continuous coding map $\pi:\Sigma_{A}\to\Omega$ with quantitative control on the exceptional set where $\pi$ fails to be injective, and identifies the topological entropy of $f|_{\Omega}$ with $\log\rho(A)$. The dimension theory of hyperbolic sets, developed by Pesin (1977, 1997), Young (1982), and Barreira et al. (1999), relies on the coding map and Markov partition constructed here. Section 10 provides a complete numerical illustration for the Smale horseshoe, computing explicit values of the adapted metric constant, stable manifold size, shadowing constant, Markov partition diameter, and grid resolution needed for rigorous symbolic coding at prescribed accuracy. Section 11 concludes the Part with a summary of the five Main Theorems and open problems. The appendix collects technical proofs and supplementary material: prerequisite results on Hölder spaces and quasi-compact operators, and standard measure-theoretic lemmas used throughout.
@@ A $C^{r}$ diffeomorphism $f:M\to M$ satisfies Axiom A if:
For an Anosov diffeomorphism, $\Omega(f)=M$ (every point is nonwandering since the stable and unstable manifolds of any point are dense). The density of periodic points follows from the Anosov Closing Lemma (Proposition 7.5), proved in Section 7. See Corollary 7.6. ∎
@@ 4.5. Quantitative Estimates
Step 3: Sharpening to rate $\lambda$. The bound in Step 2 has rate $\lambda^{\prime}=(5\lambda+3)/8$, not $\lambda$. To obtain rate $\lambda$ as stated, we use that $\lambda$ is the hyperbolicity exponent of the adapted metric, chosen with a safety factor in Subsection 3.2. Concretely, the adapted metric construction (Theorem 3.5) gives a metric $\|\cdot\|_{*}$ in which $\|Df|_{E^{s}}\|_{*}\leq\lambda$. Choosing the adapted metric with rate $\lambda_{0}<\lambda$ (any rate strictly less than $\lambda$ is allowed as long as $\lambda_{0}$ exceeds the true contraction rate of $Df|_{E^{s}}$), the same calculation as above yields rate $\lambda_{0}+C_{0}\varepsilon(1+K)$, which for $\varepsilon$ sufficiently small is at most $\lambda$. Rescaling constants accordingly proves the claim with rate exactly $\lambda$. ∎